PRELIMINARY RESULTS FROM METHYL CHLORIDE COLLISIONAL COOLING EXPERIMENTS

\underline{A. W. MANTZ}, Department of Physics and Astronomy, Connecticut College, 270 Mohegan Ave., New London, CT 06320-4196; C. BALL, and F. DE LUCIA, Department of Physics, The Ohio State University, Columbus, Ohio 43210.

We made several modifications to a low temperature cell we described last year. The modifications were intended to make the cell more robust and to provide a cell with better vacuum integrity. The changes to the cell will be described briefly. We will describe results obtained from measurements on the absorption bands of methyl chloride which are located in the seven micron spectral region.